OpenStudy (anonymous):

Kevin owes $4,645 on a credit card with a 19.7% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card? $387.08 $84.34 $429.62 $1,034.65 @Lena772 @whpalmer4 ANYONE PLEASE HELP :(

OpenStudy (whpalmer4):

Well, he's going to have to pay off at least 1/12th of the amount per month for 12 months, right, even if there's no interest. 12*84 is only about 1000, so the $84.34 answer can be safely ruled out immediately. And even at 19.7%, 12*1034 = 12,400+ so that isn't a reasonable answer either. There's your 50/50 lifeline :-)

OpenStudy (whpalmer4):

What do you know about finding the payments necessary to pay off a loan?

OpenStudy (anonymous):

i got A as the answer.. but i dont really understand it... thankyou for your help though

OpenStudy (whpalmer4):

I don't think A is correct. Starting balance is 4645. after 1 month, 4645 * 0.197/12 = 76.26 in interest is added to the balance, bringing it to 4721.26. Pay $429.62 and the balance drops to $4291.64 in the 2nd month, interest of 4291.64*0.197/12 = $70.45 is added, bringing balance to 4362.09, which then drops to 3932.47 after payment in the 3rd month, interest of 3932.47*0.197/12 = 64.56 is added, balance to 3567.41 after payment in 4th month, interest of 58.57, bal to 3196.36 in the 5th month, interest of 52.47, bal to 2819.21 in the 6th month, interest of 46.28, bal to 2435.87 in the 7th month, interest of 39.99, bal to 2046.24 in the 8th month, interest of 33.59, bal to 1650.21 in the 9th month, interest of 27.09, bal to 1247.68 in the 10th month, interest of 20.48, bal to 838.54 in the 11th month, interest of 13.77, bal to 422.69 in the 12th month, interest of 6.94, bal to 0 @sandritascandon